# WetCat: Automating Skill Assessment in Wetlab Cataract Surgery Videos

### v1

@@ 4.3. Training Settings
All neural networks are initialized with ImageNet pre-trained weights (ImageNet, ) for their respective backbones, except for ResNet3D and SAM, which are trained from scratch.
@@
The initial learning rate is set to 0.05 for ResNet3D and 0.005 for the CNN-RNN and transformer-based models. For semantic segmentation, the initial learning rate is set to 0.001 for convolutional networks and 0.0001 for SAM. For networks with pre-trained backbones, the backbone’s learning rate is set to one-tenth of the main learning rate. In all cases, the learning rate is progressively reduced during training using a polynomial decay schedule:
$lr=lr_{init}\times\left(1-\frac{iter}{total\_iter}\right)^{0.9}.$
To enhance model robustness and promote generalization, we apply a range of data augmentations during training for both tasks, including random resized cropping (scale variation between 0.8 and 1.2), random rotations (up to ±$\pm 30^{\circ}$), color jittering (brightness, contrast, and saturation adjustments up to 20%), Gaussian blurring, random sharpening, and random conversion to grayscale with a probability of 0.2.
@@
Table 4. Phase recognition performance in wetlab cataract surgery.
@@
ResNet3D 80.77 ±$\pm 1.41$ 71.17 ±$\pm 3.08$ 80.36 ±$\pm 2.59$
VGG-LSTM 85.19 ±$\pm 4.65$ 84.00 ±$\pm 4.47$ 84.85 ±$\pm 4.45$
VGG-GRU 83.90 ±$\pm 3.40$ 83.42 ±$\pm 4.14$ 84.58 ±$\pm 4.26$
VGG-BiLSTM 84.84 ±$\pm 4.72$ 85.02 ±$\pm 4.37$ 85.51 ±$\pm 4.55$
VGG-BiGRU 82.80 ±$\pm 5.13$ 82.71 ±$\pm 5.65$ 84.02 ±$\pm 5.26$
VGG-Transformer 79.75 ±$\pm 1.45$ 77.98 ±$\pm 2.77$ 80.79 ±$\pm 1.72$
@@
VGG16 UNet+ 75.55 ±$\pm 5.42$ 81.19 ±$\pm 6.37$ 78.37 85.96 ±$\pm 3.64$ 89.48 ±$\pm 3.95$ 87.72 60.00 ±$\pm 2.67$ 69.48 ±$\pm 2.63$
CPFNet 77.39 ±$\pm 5.79$ 82.21 ±$\pm 11.58$ 79.80 87.13 ±$\pm 3.81$ 89.76 ±$\pm 7.49$ 88.45 65.92 ±$\pm 3.13$ 74.72 ±$\pm 2.79$
UNetPP 75.42 ±$\pm 6.15$ 80.47 ±$\pm 8.13$ 77.94 85.84 ±$\pm 4.13$ 88.95 ±$\pm 5.17$ 87.39 60.55 ±$\pm 3.04$ 70.00 ±$\pm 2.86$
AdaptNet 72.73 ±$\pm 4.88$ 85.86 ±$\pm 10.18$ 79.29 84.12 ±$\pm 3.38$ 92.04 ±$\pm 6.29$ 88.08 65.25 ±$\pm 3.27$ 73.66 ±$\pm 3.16$
ReCal-Net 77.31 ±$\pm 5.95$ 75.52 ±$\pm 14.78$ 76.42 87.08 ±$\pm 3.92$ 85.17 ±$\pm 10.48$ 86.13 61.67 ±$\pm 2.23$ 70.55 ±$\pm 1.84$
ResNet34 CENet 77.08 ±$\pm 9.37$ 83.28 ±$\pm 11.59$ 80.18 86.72 ±$\pm 6.36$ 90.41 ±$\pm 7.32$ 88.57 48.23 ±$\pm 22.20$ 54.82 ±$\pm 25.94$
AdaptNet 74.50 ±$\pm 5.68$ 84.54 ±$\pm 12.20$ 79.52 85.27 ±$\pm 3.83$ 91.11 ±$\pm 7.69$ 88.19 68.16 ±$\pm 2.37$ 76.28 ±$\pm 2.11$
ReCal-Net 77.90 ±$\pm 5.87$ 82.79 ±$\pm 16.07$ 80.34 87.45 ±$\pm 3.84$ 89.65 ±$\pm 10.66$ 88.55 64.45 ±$\pm 1.86$ 72.55 ±$\pm 2.33$
ResNet50 UPerNet 80.41 ±$\pm 5.37$ 88.63 ±$\pm 6.41$ 84.52 89.04 ±$\pm 3.40$ 93.85 ±$\pm 3.68$ 91.44 70.81 ±$\pm 2.72$ 78.83 ±$\pm 2.31$
DeepLabV3+ 79.97 ±$\pm 6.02$ 88.66 ±$\pm 7.49$ 84.32 88.74 ±$\pm 3.86$ 93.82 ±$\pm 4.32$ 91.28 70.79 ±$\pm 2.77$ 79.01 ±$\pm 2.37$
ViT-B SAM 75.53 ±$\pm 5.00$ 91.20 ±$\pm 1.88$ 83.36 84.51 ±$\pm 4.84$ 94.97 ±$\pm 1.34$ 89.74 62.22 ±$\pm 2.81$ 74.67 ±$\pm 2.71$
SAM-LoRA 80.41 ±$\pm 3.02$ 91.63 ±$\pm 2.50$ 86.02 88.25 ±$\pm 2.48$ 94.93 ±$\pm 1.95$ 91.59 69.79 ±$\pm 2.35$ 80.62 ±$\pm 2.17$
@@
VGG16 UNet+ 44.05 ±$\pm 9.85$ 55.90 ±$\pm 10.31$
CPFNet 49.89 ±$\pm 15.92$ 59.21 ±$\pm 16.54$
UNetPP 44.42 ±$\pm 17.20$ 55.33 ±$\pm 18.23$
AdaptNet 45.17 ±$\pm 18.69$ 54.52 ±$\pm 18.51$
ReCal-Net 47.14 ±$\pm 19.96$ 56.15 ±$\pm 19.89$
ResNet34 CENet 14.53 ±$\pm 13.53$ 16.68 ±$\pm 16.86$
AdaptNet 51.38 ±$\pm 17.13$ 60.27 ±$\pm 17.18$
ReCal-Net 52.98 ±$\pm 19.68$ 60.27 ±$\pm 19.54$
ResNet50 UPerNet 68.83 ±$\pm 10.35$ 77.47 ±$\pm 9.71$
DeepLabV3+ 65.30 ±$\pm 11.76$ 74.11 ±$\pm 11.33$
ViT-B SAM 64.60 ±$\pm 4.31$ 75.14 ±$\pm 4.88$
SAM (bb) 66.88 ±$\pm 10.03$ 77.55 ±$\pm 9.41$
SAM-LoRA 74.20 ±$\pm 3.80$ 81.90 ±$\pm 3.97$
SAM-LoRA (bb) 77.57 ±$\pm 4.33$ 85.79 ±$\pm 2.97$
@@ 4.4. Experimental Results
Table 6 summarizes the rhexis segmentation performance, addressing what is arguably the most challenging task in wetlab skill assessment. The complexity stems from two factors: the limited number of annotated frames due to the short duration of the rhexis phase and the prerequisite of rhexis completion for accurate annotation, and the inherently blunt edges of the rhexis boundary.
In addition to the network architectures evaluated for anatomical structures and instruments, we also trained and assessed the Segment Anything Model (SAM) using pupil bounding box prompts, which are derived from the pupil segmentation outputs. As reported in the table, among networks operating without supervised prompts, SAM-LoRA achieved the highest performance. Notably, this performance is significantly enhanced when pupil bounding box prompts are incorporated, as reflected by the results for SAM-LoRA (bb).
@@ 5. Conclusion
In this paper, we introduce WetCat, the first curated dataset specifically designed for skill assessment in wetlab cataract surgery videos. By providing comprehensive phase annotations and semantic segmentations focused on the critical capsulorhexis and phacoemulsification phases, WetCat enables the development of interpretable, AI-driven evaluation tools. This resource addresses a critical gap in ophthalmic surgical education, fostering advancements in scalable, objective, and consistent surgical training.
@@ Acknowledgements.
This work was funded by Haag Streit AG, Switzerland. We would like to thank Finia Kowal, Alessia Bruzzo, and Ylenia Di Maro for their invaluable contributions to the meticulous annotation of the dataset.

### v3

@@ 4.3. Training Settings
All neural networks are initialized with ImageNet pre-trained weights (Deng et al., 2009) for their respective backbones, except for ResNet3D and SAM.
@@
The initial learning rate is set to 0.05 for ResNet3D and 0.005 for the CNN-RNN and transformer-based models. For semantic segmentation, the initial learning rate is set to 0.001 for convolutional networks and 0.0001 for SAM. For networks with pre-trained backbones, the backbone’s learning rate is set to one-tenth of the main learning rate. The learning rate is progressively reduced during training according to a polynomial decay schedule, as described in (Ghamsarian et al., 2024a). To enhance model robustness and promote generalization, we apply a range of data augmentations during training for both tasks, including random resized cropping (scale variation between 0.8 and 1.2), random rotations (up to $\pm 30^{\circ}$), color jittering (brightness, contrast, and saturation adjustments up to 20%), Gaussian blurring, random sharpening, and random conversion to grayscale with a probability of 0.2.
@@
Table 4. Phase recognition performance in wet-lab cataract surgery.
@@
ResNet3D 80.77 $\pm 1.41$ 71.17 $\pm 3.08$ 80.36 $\pm 2.59$
VGG-LSTM 85.19 $\pm 4.65$ 84.00 $\pm 4.47$ 84.85 $\pm 4.45$
VGG-GRU 83.90 $\pm 3.40$ 83.42 $\pm 4.14$ 84.58 $\pm 4.26$
VGG-BiLSTM 84.84 $\pm 4.72$ 85.02 $\pm 4.37$ 85.51 $\pm 4.55$
VGG-BiGRU 82.80 $\pm 5.13$ 82.71 $\pm 5.65$ 84.02 $\pm 5.26$
VGG-Transformer 79.75 $\pm 1.45$ 77.98 $\pm 2.77$ 80.79 $\pm 1.72$
@@
VGG16 UNet+ 75.55 $\pm 5.42$ 81.19 $\pm 6.37$ 78.37 85.96 $\pm 3.64$ 89.48 $\pm 3.95$ 87.72 60.00 $\pm 2.67$ 69.48 $\pm 2.63$
CPFNet 77.39 $\pm 5.79$ 82.21 $\pm 11.58$ 79.80 87.13 $\pm 3.81$ 89.76 $\pm 7.49$ 88.45 65.92 $\pm 3.13$ 74.72 $\pm 2.79$
UNetPP 75.42 $\pm 6.15$ 80.47 $\pm 8.13$ 77.94 85.84 $\pm 4.13$ 88.95 $\pm 5.17$ 87.39 60.55 $\pm 3.04$ 70.00 $\pm 2.86$
AdaptNet 72.73 $\pm 4.88$ 85.86 $\pm 10.18$ 79.29 84.12 $\pm 3.38$ 92.04 $\pm 6.29$ 88.08 65.25 $\pm 3.27$ 73.66 $\pm 3.16$
ReCal-Net 77.31 $\pm 5.95$ 75.52 $\pm 14.78$ 76.42 87.08 $\pm 3.92$ 85.17 $\pm 10.48$ 86.13 61.67 $\pm 2.23$ 70.55 $\pm 1.84$
ResNet34 CENet 77.08 $\pm 9.37$ 83.28 $\pm 11.59$ 80.18 86.72 $\pm 6.36$ 90.41 $\pm 7.32$ 88.57 48.23 $\pm 22.20$ 54.82 $\pm 25.94$
AdaptNet 74.50 $\pm 5.68$ 84.54 $\pm 12.20$ 79.52 85.27 $\pm 3.83$ 91.11 $\pm 7.69$ 88.19 68.16 $\pm 2.37$ 76.28 $\pm 2.11$
ReCal-Net 77.90 $\pm 5.87$ 82.79 $\pm 16.07$ 80.34 87.45 $\pm 3.84$ 89.65 $\pm 10.66$ 88.55 64.45 $\pm 1.86$ 72.55 $\pm 2.33$
ResNet50 UPerNet 80.41 $\pm 5.37$ 88.63 $\pm 6.41$ 84.52 89.04 $\pm 3.40$ 93.85 $\pm 3.68$ 91.44 70.81 $\pm 2.72$ 78.83 $\pm 2.31$
DeepLabV3+ 79.97 $\pm 6.02$ 88.66 $\pm 7.49$ 84.32 88.74 $\pm 3.86$ 93.82 $\pm 4.32$ 91.28 70.79 $\pm 2.77$ 79.01 $\pm 2.37$
ViT-B SAM 75.53 $\pm 5.00$ 91.20 $\pm 1.88$ 83.36 84.51 $\pm 4.84$ 94.97 $\pm 1.34$ 89.74 62.22 $\pm 2.81$ 74.67 $\pm 2.71$
SAM-LoRA 80.41 $\pm 3.02$ 91.63 $\pm 2.50$ 86.02 88.25 $\pm 2.48$ 94.93 $\pm 1.95$ 91.59 69.79 $\pm 2.35$ 80.62 $\pm 2.17$
@@
VGG16 UNet+ 44.05 $\pm 9.85$ 55.90 $\pm 10.31$
CPFNet 49.89 $\pm 15.92$ 59.21 $\pm 16.54$
UNetPP 44.42 $\pm 17.20$ 55.33 $\pm 18.23$
AdaptNet 45.17 $\pm 18.69$ 54.52 $\pm 18.51$
ReCal-Net 47.14 $\pm 19.96$ 56.15 $\pm 19.89$
ResNet34 CENet 14.53 $\pm 13.53$ 16.68 $\pm 16.86$
AdaptNet 51.38 $\pm 17.13$ 60.27 $\pm 17.18$
ReCal-Net 52.98 $\pm 19.68$ 60.27 $\pm 19.54$
ResNet50 UPerNet 68.83 $\pm 10.35$ 77.47 $\pm 9.71$
DeepLabV3+ 65.30 $\pm 11.76$ 74.11 $\pm 11.33$
ViT-B SAM 64.60 $\pm 4.31$ 75.14 $\pm 4.88$
SAM (bb) 66.88 $\pm 10.03$ 77.55 $\pm 9.41$
SAM-LoRA 74.20 $\pm 3.80$ 81.90 $\pm 3.97$
SAM-LoRA (bb) 77.57 $\pm 4.33$ 85.79 $\pm 2.97$
@@ 4.4. Experimental Results
Table 6 summarizes the rhexis segmentation performance, addressing what is arguably the most challenging task in wet-lab skill assessment. The complexity stems from two factors: the limited number of annotated frames due to the short duration of the rhexis, and the inherently blunt edges of the rhexis boundary. For this task, in addition to the network architectures evaluated for anatomical structures and instruments, we trained and evaluated the Segment Anything Model (SAM) using pupil bounding box prompts derived from the pupil segmentation results. As reported in the table, among networks operating without supervised prompts, SAM-LoRA achieved the highest performance. Notably, this performance is significantly enhanced when pupil bounding box prompts are incorporated, as reflected by the results for SAM-LoRA (bb).
@@ 5. Conclusion
In this paper, we introduce WetCat, the first curated dataset specifically designed for skill assessment in wet-lab cataract surgery videos. By providing comprehensive phase annotations and semantic segmentations focused on the critical capsulorhexis and phacoemulsification phases, WetCat enables the development of interpretable, AI-driven evaluation tools. This resource addresses a critical gap in ophthalmic surgical education, fostering advancements in scalable, objective, and consistent surgical training.
@@ Acknowledgements.
This work was funded by HaagStreit AG. We would like to thank Finia Kowal, Alessia Bruzzo, and Ylenia Di Maro for their invaluable contributions to the meticulous annotation of the dataset.